Two aqueous solutions denoted S1 and S2 are prepared in the laboratory, solution S1 contains 37g of copper sulfate dissolved in 100 ml of water and solution S2 contains 50 g of the same solute dissolved in 200 ml of water.


1- Show that the concentration (C) of S1 is 370 g\l and that of S2 is 250 g L.


2- Specify which of the two solutions is the most concentrated and which is the most dilute.


3- Explain how the S1 solution can be made more dilute.


4- Suggest a method to increase the concentration of S2.


1
Expert's answer
2021-04-11T23:51:53-0400

1) С = m / V

C1 = 37 g / 0.1 L = 370 g/L

C2 = 50 g / 0.2 L = 250 g/L


2) S1 is more concentrated solution, than S2. Respectively, S2 is more diluted solution than S1.


3) S1 solution can be made more dilute by adding more solvent or removing solute.


4) The concentration of S2 can be increased through adding some more sollute or removing part of the solvent.